Edward C. answered 12/01/15
Tutor
5.0
(439)
Caltech Grad for math tutoring: Algebra through Calculus
Sometimes it's easier to calculate the probability that something will not happen than the probability that it will happen. Since these 2 events are mutually exclusive and span the entire probability space their probabilities must sum to 1, so P(have to be replaced) = 1 - P(not have to be replaced). So see if you can figure out P(not have to be replaced).
Hint: Since P(bad egg) = 0.05, what is P(good egg)?
What is P(all 12 eggs in a dozen good)?
If you need more help comment back